If a clock strikes every hour the same amount of times as that hour, how many strikes in 24hours.
Answer by RAY100(1637)

This is basically an arithmetic series 1,2,3,,,24
.
form is a(n) = a(1) +(n-1)d,,,,,where d is difference btwn terms, and n = 3 terms
.
a(24) = 1+(24-1)1 = 24
.
The sum of an arithmetic series is,,S= n{ a(1) +a(n)} /2
.
S= 24(1+24)/2 = 300
.
.
check,,,use calc to manually add series,,,300,,,,ok